# Message-Passing Monte Carlo: Generating low-discrepancy point sets via Graph Neural Networks

### v1

@@ 3.3 Training objective
Our ultimate goal is to minimize the star-discrepancy $D^{*}$ (1). However, $D^{*}$ cannot serve as the training objective, as (i) $D^{*}$ is computationally infeasible to calculate for high dimensions $d$ and large number of points $N$ (it has been shown to be an NP-hard problem in (Gnewuch et al., 2009) and in fact, it is even W$[1]$-hard in $d$ (Giannopoulos et al., 2012)); (ii) the training objective should not only be computationally feasible but rather very efficient to compute, as it needs to be evaluated at every step of the training procedure (i.e., for every step of the gradient descent method) resulting in potentially thousands of evaluations to train only a single model, and (iii) the training objective needs to be differentiable in order to be used in the context of gradient-based learning. It turns out, we can derive a training objective resolving all three issues while simultaneously minimizing $D^{*}$ by leveraging previous work on the ${\mathcal{L}}_{p}$-discrepancy,
@@
Clearly, the star-discrepancy $D^{*}$ can be derived as a special case of (LABEL:eq:Lp_disc) with $p=\infty$. Here, we focus on the case of $p=2$ as our training objective, since instead of computing the integral in (LABEL:eq:Lp_disc), we can leverage its closed-form expression, known as Warnock’s formula (Warnock, 1972),
@@ 3.4 Extension to higher dimensions
In many practical problems, particularly in engineering and finance, the dimension $d$ of the problem can be very large. This necessitates extending low-discrepancy sequences to the high dimensional case of $d\gg 1$. However, it is known (Morokoff & Caflisch, 1994; Wang & Sloan, 2008) that the ${\mathcal{L}}_{2}$-discrepancy fails to identify superior distributional properties of low-discrepancy point sets over random samples as the dimension increases. Indeed, in high dimensions the classical ${\mathcal{L}}_{2}$-discrepancy of low-discrepancy point sets behaves like $\mathcal{O}(1/\sqrt{N})$, the same as for random points, for moderate values of $N$, while an improved order close to $\mathcal{O}(1/N)$ can only be seen for extremely large $N$. Empirical evidence for these last claims can be found in the discrepancy plots contained in (Morokoff & Caflisch, 1994).
To this end, we suggest to base our new training objective for higher-dimensional generation of low-discrepancy points on the Hickernell ${\mathcal{L}}_{p}$-discrepancy (Hickernell, 1998),
@@
where $\emptyset\neq s_{k}\sim\mathcal{P}(\{1,\dots,d\})$ are randomly sampled subsets of coordinate indices for each $k=1,\dots,K$, thus requiring to compute the ${\mathcal{L}}_{p}$-discrepancy only $K$ times.
@@
As a further advantage to this framework, we highlight its inherent flexibility. Specifically, employing the modified Hickernell discrepancy as the training objective represents a first step towards an adaptive QMC sampling method tailored for specific problems. It is widely recognized that for many problems, the effective dimension—essentially, the number of dimensions capturing the majority of the problem’s variability—is often significantly lower than the nominal dimension; for full details, refer to (Caflisch et al., 1997). Therefore, during high-dimensional training, prioritizing sampling from specific lower dimensional projections will yield a $d$-dimensional point set that is highly uniformly distributed in those same projections identified during training. This approach effectively creates a custom-made point set, optimized for problems that primarily depend upon particular dimensions.
@@ 4 Empirical results
In this section, we present empirical results comparing MPMC points to current state-of-the-art low-discrepancy point sets. More concretely, we demonstrate superior distributional properties of MPMC over other low-discrepancy point sets mainly with respect to the star-discrepancy $D^{*}$.
As described above, computing the star-discrepancy is usually a difficult task. Being a discrete problem, depending on the number of points, a crude search for the worst distribution over all admissible test boxes is often extremely slow in low dimensions, and intractable even in dimension as small as four or five. The best known exact algorithm to calculate the star-discrepancy is the so-called Dopkin, Eppstein and Mitchell (DEM) algorithm (Dobkin et al., 1996) which runs in $\mathcal{O}(N^{1+d/2})$ time. In all experiments, the results of which are presented shortly, to calculate the star discrepancy we either use a simple crude search or a parallelized version of the DEM algorithm from (Clément et al., 2023) to speed up calculation when necessary. The interested reader is recommended to consult (Doerr et al., 2014) and references therein for more information on the calculation of the star-discrepancy, and indeed the computation of other discrepancy measures.
@@ 4.1 Low-dimensional generation of MPMC points
Here, we focus on generating MPMC points within a lower-dimensional setting particularly because this area has recently attracted significant attention (Clément et al., 2022, 2024, 2023) providing a solid basis for comparison. Additionally, QMC methods are known to perform very well on many complex applications as long as the low-dimensional projections of the underlying point set have low discrepancy. Therefore, generating 2- or 3-dimensional point sets with strong distributional properties are of paramount interest toward building competitive constructions performing well on a variety of problems.
Figure 4: Star-discrepancy $D^{*}$ of Halton, Sobol’, Subset Selection, Hammersley, Fibonacci, and MPMC for increasing number of points $N=20,\dots,500$ in $d=2$.
We compare the irregularity of our MPMC point sets with a truncation to the first $N$ points of the widely used Sobol’ and Halton sequences. We note that MPMC points are sets optimized for $N$ chosen in advance, whereas a key advantage of the Sobol’ and Halton sequences are that they are built for repeated sampling and retain a low-discrepancy for all values of $N$. Therefore, in addition, we provide comparison with state-of-the-art point sets derived from the subset selection method from (Clément et al., 2022, 2024), the Hammersley construction in base $1+\sqrt{2}$ as introduced in (Kirk et al., 2023) and the Fibonacci set defined as $\left\{(i/n,\{i\varphi\}):i\in\{0,\ldots,n-1\}\right\}$ where $\varphi$ represents the golden ratio, the notation $\{x\}$ denotes the fractional part of $x\in\mathbb{R}$. All three of these sets are recognized for having among the lowest star-discrepancy for given $N$ in two dimensions. Fig. 4 shows the star-discrepancy of MPMC, Sobol’, Halton, subset selection, Hammersley and Fibonacci sets in two dimensions for increasing number of points $N=20,\dots,500$. We can see that MPMC significantly outperforms all other methods with respect to the star-discrepancy. In fact, the star-discrepancy of MPMC is on average $1.5$ times smaller than that of the current state-of-the-art Fibonacci construction, and on average more than $2.5$ times smaller than Sobol’ or Halton points. We provide the exact values of Fig. 4 in the Supporting Information (SI) such that it can serve as a benchmark for future methods.
@@ 4.2 Optimality of MPMC point sets
Much of the past research on low-discrepancy point sets focused on achieving star-discrepancy with optimal asymptotic order in $N$ for implementation in quasi-Monte Carlo methods. However, there has been a recent surge in interest in finding point sets that minimize discrepancy for fixed $N$ and $d$. The main contribution in this direction was given in (Clément et al., 2023), where the authors constructed optimal star-discrepancy point sets in two and three dimensions. Naturally, we are interested in comparing MPMC points to these optimal formulations. The results of the optimal star-discrepancy comparison in two dimensions are presented in Table 1 and the three dimensional case is found in the SI.
@@
N 1 2 3 4 5
Fibonacci 1.0 0.6909 0.5880 0.4910 0.3528
Optimal $1/\varphi$ 0.366 0.2847 0.25 0.2
MPMC $1/\varphi$ 0.366 0.2847 0.25 0.2
N 6 7 8 9 10
Fibonacci 0.3183 0.2728 0.2553 0.2270 0.2042
Optimal 0.1667 0.15 0.1328 0.1235 0.1111
MPMC 0.1692 0.1508 0.1354 0.1240 0.1124
N 11 12 13 14 15
Fibonacci 0.1857 0.1702 0.1571 0.1459 0.1390
Optimal 0.1030 0.0952 0.0889 0.0837 0.0782
MPMC 0.1058 0.0975 0.0908 0.0853 0.0794
N 16 17 18 19 20
Fibonacci 0.1486 0.1398 0.1320 0.1251 0.1188
Optimal 0.0739 0.0699 0.0666 0.0634 0.0604
MPMC 0.0768 0.0731 0.0699 0.0668 0.0640
We can see that the star-discrepancy of MPMC points is very close to the star-discrepancy of the optimal point sets, and in fact match it exactly for small $N$. Moreover, the star-discrepancy of the Fibonacci set is far off the optimal values, i.e, approximately by a factor of $2$. Finally, it is worth highlighting that the computation of the optimal points requires to solve a non-linear programming problem and takes approximately $18$ days to compute. In contrast to that, our model generating MPMC points can be trained from scratch in less than $5$ minutes on a conventional GPU machine.
@@ 4.3 MPMC generation in high dimensions
As discussed in Section 3.4, the efficacy of discrepancy measures to justly evaluate irregularity of distribution is flawed in higher dimensions. Therefore, to alternatively assess the quality of the distribution of higher dimensional point sets and sequences, motivated by the Koksma-Hlawka inequality (2), we will implement high dimensional MPMC points in an integral arising in a real-world problem from computational finance previously studied in (Lemieux & Owen, 2002; Wang & Sloan, 2005; Faure & Lemieux, 2009).
The primary goal is to accurately estimate the value at time 00 of an Asian call option on an underlying asset that follows a log-normal distribution. Complete details of the problem formulation are provided in the SI. Table 2 shows the absolute errors observed when implementing Halton, Sobol’ and MPMC constructions. With our chosen parameters, this problem is known (Lemieux & Owen, 2002) to exhibit more than 97% of its variability in dimensions one, two, and three. Therefore, we train a 32-dimensional MPMC point set while emphasizing the $1$-$3$-dimensional projections as described in Section 3.4. We report the average absolute error of an MPMC training batch, which is selected based on the minimal Hickernell ${\mathcal{L}}_{2}$-discrepancy restricted to $1$-$3$-dimensional projections.
Table 2: Approximation error of an Asian call option pricing of MPMC, Halton and Sobol’.
N 32 64 128 256 512
Halton 6.4566 4.1580 3.5590 2.8046 1.9376
Sobol 1.2163 1.3545 0.9470 0.6043 0.4784
MPMC 1.3836 0.8129 0.4937 0.2317 0.1019
Once again, the 32-dimensional MPMC point sets show significant enhancements over Halton and Sobol’. This improvement is particularly notable at higher values of $N$, where MPMC outperforms Sobol’ by a factor of approximately $5$, and Halton by a factor of approximately $19$. The efficiency observed in this high-dimensional problem is promising and suggests a superior uniformity of MPMC points. In fact, the observed improvements may partially be explained by the uniformity held in lower-dimensional projections when compared to the traditional choices of QMC point sets. We refer to the SI for further discussion.
@@ 4.4 Ablations
How does the graph structure influence the performance? To answer this, we compute the average ${\mathcal{L}}_{2}$-discrepancy of several trained MPMC models for increasing values of the nearest neighbor radius $r$ in (5) ranging from 00 to $1$ for different number of points $N=64,128,1024$, and plot the results in the SI. These results lead to two important observations. First, not using a graph structure at all, i.e., setting $r=0$ resulting in Deepsets (Zaheer et al., 2017), significantly impairs the performance of MPMC, reaching average ${\mathcal{L}}_{2}$-discrepancy values that are $9$ to $40$ times worse than using a graph structure. The second observation is that although the performance of MPMC is relatively stable for any choice of $r>0$, including the radius $r$ in hyperparameter tuning can help achieve point sets with minimal discrepancy.
What is the role of the GNN architecture used in MPMC? While we base MPMC on message-passing neural networks (MPNNs) (Gilmer et al., 2017), other GNN architectures such as Graph Convolutional Networks (GCNs) (Kipf & Welling, 2017), or Graph Attention Networks (GATs) can be used in this context as well. To check this, we train MPMC based on MPNNs, GCNs, and GATs for three different number of points $N=64,256,1024$ and show the ${\mathcal{L}}_{2}$-discrepancy in the SI. Based on these results, we conclude that GCNs and GATs outperform each other based on the number of points $N$. At the same time, MPNNs consistently yield point sets with the lowest discrepancy values among all three considered GNN architectures.
@@ 5 Discussion
Low-discrepancy points play a central role in many applications in science and engineering. In this article, we have proposed MPMC, the first machine learning approach to generate new sets of low-discrepancy points. Inspired by the geometric nature of constructing such point sets, we base our MPMC approach on GNNs. Choosing an adequate training objective, i.e., closed-form solution of the ${\mathcal{L}}_{2}$-discrepancy, we show that MPMC successfully transforms (random) input points into point sets with low discrepancy. Moreover, we extend this framework to higher dimensions, by training with an approximation of the Hickernell ${\mathcal{L}}_{2}$-discrepancy. We further present an extensive empirical evaluation to illustrate different aspects of the proposed MPMC approach, highlighting the superior distributional properties of MPMC points compared to previous state-of-the-art methods. Finally, we carefully ablate key components of our MPMC model, yielding deeper empirical insights.
MPMC represents a novel and efficient way of generating point sets with very low discrepancy. In fact, MPMC is shown to obtain optimal or near-optimal discrepancy in near-real time. This is crucial for computationally expensive applications, where MPMC will lead to potentially significantly lower absolute errors compared to previous methods. Moreover, the generality of the MPMC framework allows for designing tailor-made QMC points that exploit specific structures of the problem at hand.
The aim of this paper was to generate point sets with low discrepancy for a fixed dimension and fixed number of points. On the other hand, many important applications require repeated sampling resulting in low-discrepancy sequences and not fixed point sets. Thus, one important aspect of future work will be to extend our MPMC point sets to MPMC sequences. Another essential part of this paper was on studying the distributional properties of MPMC points, in contrast to simply showcasing their performance in practical applications. However, based on the empirical evidence in this paper, we expect MPMC point sets to excel in various applications. Motivated by this, we would like to apply MPMC to various problems in science and engineering as future work.
Acknowledgements

### v2

@@ 2.3 Training objective
Our ultimate goal is to minimize the star-discrepancy $D^{*}$ (1). However, $D^{*}$ cannot serve as the training objective, as (i) $D^{*}$ is computationally infeasible to calculate for high dimensions $d$ and large number of points $N$, (ii) the training objective should not only be computationally feasible but rather very efficient to compute, as it needs to be evaluated at every step of the training procedure (i.e., for every step of the gradient descent method) resulting in potentially thousands of evaluations to train only a single model, and (iii) the training objective needs to be sufficiently differentiable (i.e., that can be handled by automatic differentiation packages such as (46)) in order to be used in the context of gradient-based learning. It turns out, we can derive a training objective resolving all three issues while simultaneously reducing $D^{*}$ by leveraging previous work on the ${\mathcal{L}}_{p}$-discrepancy,
@@
Clearly, the star-discrepancy $D^{*}$ can be derived as a special case of (LABEL:eq:Lp_disc) with $p=\infty$. Here, we focus on the case of $p=2$ as our training objective, since instead of computing the integral in (LABEL:eq:Lp_disc), we can leverage its closed-form expression, known as Warnock’s formula (47),
@@ 2.4 Extension to higher dimensions
In many practical problems, particularly in engineering and finance, the dimension $d$ of the problem can be very large. This necessitates extending low-discrepancy sequences to the high dimensional case of $d\gg 1$. However, it is known (48, 49) that the ${\mathcal{L}}_{2}$-discrepancy fails to identify superior distributional properties of low-discrepancy point sets over random samples as the dimension increases. Indeed, in high dimensions the classical ${\mathcal{L}}_{2}$-discrepancy of low-discrepancy point sets behaves like $\mathcal{O}(1/\sqrt{N})$, the same as for random points, for moderate values of $N$, while an improved order close to $\mathcal{O}(1/N)$ can only be seen for extremely large $N$. Empirical evidence for these last claims can be found in the discrepancy plots contained in (48).
To this end, we suggest to base our new training objective for higher-dimensional generation of low-discrepancy points on the Hickernell ${\mathcal{L}}_{p}$-discrepancy (50),
@@
where $\emptyset\neq s_{k}\sim\mathcal{P}(\{1,\dots,d\})$ are randomly sampled subsets of coordinate indices for each $k=1,\dots,K$, thus requiring to compute the ${\mathcal{L}}_{p}$-discrepancy only $K$ times. More specifically, we sample $s_{k}$ by first selecting the dimension $d_{k}$ of the projection uniformly at random, and then sampling the projection itself uniformly at random from the set of all $d_{k}$-dimensional projections. This method ensures that the dimensions of the projections are uniformly distributed during training. Alternatively, we could randomly sample the dimensions of the projections according to a binomial distribution.
@@
As a further advantage to this framework, we highlight its inherent flexibility. Specifically, employing the modified Hickernell discrepancy as the training objective represents a first step towards an adaptive QMC sampling method tailored for specific problems. It is widely recognized that for many problems, the effective dimension—essentially, the number of dimensions capturing the majority of the problem’s variability—is often significantly lower than the nominal dimension; for full details, refer to (51). Therefore, during high-dimensional training, assuming that the important subsets of variables are known or identified in advance, e.g., by functional ANOVA methods (52), prioritizing sampling from specific lower dimensional projections will yield a $d$-dimensional point set that is highly uniformly distributed in those same projections identified during training. This approach effectively creates a custom-made point set, optimized for problems that primarily depend upon particular subsets of variables.
@@ 3 Empirical results
In this section, we present empirical results comparing MPMC points to current state-of-the-art low-discrepancy point sets. More concretely, we demonstrate superior distributional properties of MPMC over other low-discrepancy point sets mainly with respect to the star-discrepancy, $D^{*}$.
As previously mentioned, computing the star-discrepancy is typically a challenging task. The DEM algorithm (53) denotes the fastest method to compute the exact star-discrepancy with a significantly reduced complexity, running in $\mathcal{O}(N^{1+d/2})$ time. In all experiments, the results of which are presented shortly, to calculate the star discrepancy we either use a simple crude search or a parallelized version of the DEM algorithm from (54) to speed up calculation when necessary. The interested reader is recommended to consult (55) and references therein for more information on the calculation of the star-discrepancy, and indeed the computation of other discrepancy measures.
@@ 3.1 Low-dimensional generation of MPMC points
Here, we focus on generating MPMC points within a lower-dimensional setting particularly because this area has recently attracted significant attention (27, 28, 29) providing a solid basis for comparison.
Figure 4: Star-discrepancy $D^{*}$ of Halton, Sobol’, lifted Sobol’, Subset Selection, Hammersley, Fibonacci, and MPMC for increasing number of points $N=20,\dots,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1020}$ in $d=2$.
We compare the irregularity of our MPMC point sets with a truncation to the first $N$ points of the widely used Sobol’ and Halton sequences. We note that MPMC points are sets optimized for $N$ chosen in advance, whereas a key advantage of the Sobol’ and Halton sequences are that they are built for repeated sampling and retain a low-discrepancy for all values of $N$. Therefore, in addition, we provide comparison with state-of-the-art point sets derived from the subset selection method from (27, 28), lifted Sobol’ (the first $N$ terms of a one-dimensional Sobol’ sequence concatenated with $i/N$ for $i\in\{0,\dots,N-1\}$), the Hammersley construction in base $1+\sqrt{2}$ as introduced in (56) and the Fibonacci set defined as $\left\{(i/N,\{i\varphi\}):i\in\{0,\ldots,N-1\}\right\}$ where $\varphi$ represents the golden ratio, the notation $\{x\}$ denotes the fractional part of $x\in\mathbb{R}$. All four of these sets are recognized for having among the lowest star-discrepancy for given $N$ in two dimensions. Fig. 4 shows the star-discrepancy of MPMC, Sobol’, Halton, subset selection, lifted Sobol’, Hammersley and Fibonacci sets in two dimensions for increasing number of points $N=20,\dots,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1020}$. We can see that MPMC significantly outperforms all other methods with respect to the star-discrepancy. In fact, the star-discrepancy of MPMC is on average $1.3$ times smaller than that of the current state-of-the-art Fibonacci construction, and on average more than $2.2$ times smaller than Sobol’ or Halton points. We provide the exact values of Fig. 4 in the Supporting Information (SI) such that it can serve as a benchmark for future methods. We further provide the ${\mathcal{L}}_{2}$-discrepancy values of MPMC together with a discussion outlining their relevance in the SI.
@@ 3.2 Optimality of MPMC point sets
Much of the past research on low-discrepancy point sets focused on achieving star-discrepancy with optimal asymptotic order in $N$ for implementation in quasi-Monte Carlo methods. However, there has been a recent surge in interest in finding point sets that minimize discrepancy for fixed $N$ and $d$. The main contribution in this direction was given in (29), where the authors constructed optimal star-discrepancy point sets in two and three dimensions. Naturally, we are interested in comparing MPMC points to these optimal formulations. The results of the optimal star-discrepancy comparison in two dimensions are presented in Table 1 and the three dimensional case is found in the SI.
@@
N 1 2 3 4 5 6 7
Fibonacci 1.0 0.6909 0.5880 0.4910 0.3528 0.3183 0.2728
Optimal $1/\varphi$ 0.3660 0.2847 0.2500 0.2000 0.1667 0.1500
MPMC $1/\varphi$ 0.3660 0.2847 0.2500 0.2000 0.1692 0.1508
N 8 9 10 11 12 13 14
Fibonacci 0.2553 0.2270 0.2042 0.1857 0.1702 0.1571 0.1459
Optimal 0.1328 0.1235 0.1111 0.1030 0.0952 0.0889 0.0837
MPMC 0.1354 0.1240 0.1124 0.1058 0.0975 0.0908 0.0853
N 15 16 17 18 19 20 21
Fibonacci 0.1390 0.1486 0.1398 0.1320 0.1251 0.1188 0.1131
Optimal 0.0782 0.0739 0.0699 0.0666 0.0634 0.0604 0.0580
MPMC 0.0794 0.0768 0.0731 0.0699 0.0668 0.0640 0.0606
We can see that the star-discrepancy of MPMC points is very close to the star-discrepancy of the optimal point sets, and in fact match it exactly for small $N$. Moreover, the star-discrepancy of the Fibonacci set is far off the optimal values, i.e., approximately by a factor of $2$. Finally, it is worth highlighting, as reported in (29), that the computation of the optimal points requires to solve a non-linear programming problem and takes approximately $18$ days to compute for the case of $d=3$ and $N=8$. In contrast to that, MPMC was trained from scratch in $72$ seconds on an NVIDIA GeForce RTX 2080 Ti GPU for the same case.
@@ 3.3 MPMC generation in high dimensions
As discussed in Section 2.4, the efficacy of discrepancy measures to justly evaluate irregularity of distribution is flawed in higher dimensions. Therefore, to alternatively assess the quality of the distribution of higher dimensional point sets and sequences, motivated by the Koksma-Hlawka inequality (2), we will implement high dimensional MPMC points in an integral arising in a real-world problem from computational finance previously studied in (52, 57, 58).
The primary goal is to accurately estimate the value at time 00 of an Asian call option on an underlying asset that follows a log-normal distribution. Complete details of the problem formulation are provided in the SI. With our chosen parameters, this problem is known (52) to exhibit more than 97% of its variability in dimensions one, two, and three. Table 2 shows the absolute errors observed when implementing Hammersley, lattice, Sobol’ and MPMC constructions. We train a 32-dimensional MPMC point set while emphasizing the $1$-$3$-dimensional projections as described in Section 2.4. Likewise, we utilize the custom QMC software LatNet Builder (59) to construct a rank-1 lattice by the component-by-component construction (60, 61) placing importance on the $1$-$3$-dimensional projections. We report the average absolute error of an MPMC training batch, which is selected based on the minimal Hickernell ${\mathcal{L}}_{2}$-discrepancy restricted to $1$-$3$-dimensional projections.
Table 2: Approximation error of an Asian call option pricing of MPMC, Hammersley, a rank-1 lattice, and Sobol’.
N 32 64 128 256 512 1024
Hammersley 6.449 4.125 3.575 2.817 1.947 1.296
Rank-1 lattice 5.636 4.638 1.331 2.151 0.180 0.203
Sobol’ 1.235 1.373 0.965 0.623 0.497 0.290
MPMC 1.402 0.831 0.512 0.250 0.120 0.055
The 32-dimensional MPMC point sets show significant enhancements over previous methods. This improvement is particularly notable at higher values of $N$, where MPMC outperforms rank-1 lattice by a factor of approximately $4$, Sobol’ by a factor of $5$, and Hammersley by a factor of $24$. A particularly promising feature is the gain shown by MPMC points through targeted training, surpassing LatNet Builder’s targeted rank-1 lattice. The efficiency observed in this high-dimensional problem suggests a superior uniformity of MPMC points. In fact, the observed improvements may partially be explained by the uniformity held in lower-dimensional projections when compared to the traditional choices of QMC point sets. We refer to the SI for further discussion.
@@ 3.4 Ablations
How does the graph structure influence the performance? To answer this, we compute the average ${\mathcal{L}}_{2}$-discrepancy of several trained MPMC models for increasing values of the nearest neighbor radius $r$ in (4) ranging from 00 to $\sqrt{d}$ for different number of points $N=64,128,1024$, and plot the results in the SI. These results lead to two important observations. First, not using a graph structure at all, i.e., setting $r=0$ resulting in Deepsets (62), significantly impairs the performance of MPMC, reaching average ${\mathcal{L}}_{2}$-discrepancy values that are $9$ to $40$ times worse than using a graph structure. The second observation is that although the performance of MPMC is relatively stable for any choice of $r>0$, including the radius $r$ in hyperparameter tuning can help achieve point sets with minimal discrepancy.
What is the role of the GNN architecture used in MPMC? While we base MPMC on message-passing neural networks (MPNNs) (42), other GNN architectures such as Graph Convolutional Networks (GCNs) (38), or Graph Attention Networks (GATs) can be used in this context as well. To check this, we train MPMC based on MPNNs, GCNs, and GATs for three different number of points $N=64,256,1024$ and show the ${\mathcal{L}}_{2}$-discrepancy in the SI. Based on these results, we conclude that GCNs and GATs outperform each other based on the number of points $N$. At the same time, MPNNs consistently yield point sets with the lowest discrepancy values among all three considered GNN architectures.
@@ 4 Discussion
Low-discrepancy points play a central role in many applications in science and engineering. In this article, we have proposed MPMC, the first machine learning approach to generate new sets of low-discrepancy points. Inspired by the geometric nature of constructing such point sets, we base our MPMC approach on GNNs. Choosing an adequate training objective, i.e., closed-form solution of the ${\mathcal{L}}_{2}$-discrepancy, we show that MPMC successfully transforms (random) input points into point sets with low discrepancy. Moreover, we extend this framework to higher dimensions, by training with an approximation of the Hickernell ${\mathcal{L}}_{2}$-discrepancy. We further present an extensive empirical evaluation to illustrate different aspects of the proposed MPMC approach, highlighting the superior uniformity properties of MPMC points compared to previous state-of-the-art methods. Finally, we carefully ablate key components of our MPMC model, yielding deeper empirical insights.
MPMC represents a novel and efficient way of generating point sets with very low discrepancy. In fact, MPMC is empirically shown to obtain optimal or near-optimal discrepancy for every dimension and the number of points for which the optimal discrepancy can be determined. This is crucial for computationally expensive applications, where MPMC will lead to potentially significantly lower absolute errors compared to previous methods. Moreover, the generality of the MPMC framework allows for designing tailor-made QMC points that exploit specific structures of the problem at hand.
The aim of this paper was to generate point sets with low discrepancy for a fixed dimension and fixed number of points. On the other hand, many important applications require repeated sampling resulting in low-discrepancy sequences and not fixed point sets. Thus, one important aspect of future work will be to extend our MPMC point sets to MPMC sequences. Lastly, based on the superior discrepancy performance, we expect MPMC point sets to excel in various applications. Motivated by this, we would like to apply MPMC to various problems in science and engineering as future work.
[…]